Question regarding about Quadratic Function.
This is the problem:
A quadratic function is given below:
f(x)=3x^2-x+6
What is f(2)?
A. 40
B. 28
C. 16
D. 4
If you may explain how to solve this I would really appreciate the help. I don`t know any strategies that can help me solve this problem aside from having a calculator.
Thank you SO much for the time and effort.
Regards,
Alex H.

Answer by nerdybill(7384)

A quadratic function is given below:
f(x)=3x^2-x+6
What is f(2)?
A. 40
B. 28
C. 16
D. 4
.
f(x)=3x^2-x+6
f(x) is a "function of x" -- meaning for a particular value of x you can find the value of f(x).
In your case, when they say "what is f(2)"? They are asking you what the value of f(x) when x=2. So, simply plug in the value of 2 for x and solve:
f(x)=3x^2-x+6
f(2)=3(2^2)-2+6
f(2)=3(4)-2+6
f(2)=12-2+6
f(2)=10+6
f(2)=16